OpenStudy (anonymous):

Find the area of the region bounded by the functions f(x) = x4 and g(x) = |x|.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

since both are even functions, you can focus on the portion in Q1 then just double that result to get the full area

jimthompson5910 (jim_thompson5910):

in Q1, x > 0 when x > 0, |x| = x So when x > 0, g(x) = x

jimthompson5910 (jim_thompson5910):

solve x^4 = x to find out where the two functions cross

OpenStudy (anonymous):

x=1 and x=0

OpenStudy (irishboy123):

\[\checkmark\]

jimthompson5910 (jim_thompson5910):

so you need to compute \[\Large 2*\int_{0}^{1}\left(g(x) - f(x)\right)dx\] where f(x) = x^4 g(x) = x

OpenStudy (loser66):

I don't think we get 2 in the front @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes because of symmetry about the y axis

OpenStudy (loser66):